Tech Staff Meeting Notes January 6, 2015

The top priority is always repair of down stations.
NM problem stations
ET problem stations

Duty Table updated June 2014


News & Highlights

Due to power and lack of sunlight issues, station outages were not updated for this meeting--they are the same as the December meeting.

We've been having an unusually overcast winter so far and Mitch initiated a discussion on the anss-netops list. There were a few good suggestions but basically it boils down to this being a relatively rare climatic event so herculean efforts may not be justified. PWLA will get a bigger solar panels. The AG stations will get AC (in addition to existing solar) and we need to find out why the trillium at LNXT is drawing 100mw more than it should.

rtpm on hector is very slow. This is most likely due to the REQ3 and REQ3a processes running on hector (produces this and this We should move rtpm and its associated rtpd to another machine. rtp2rtpd on achilles will need to be modified to connect with the new machine.

RTPD on starworm failed. That's twice. Need to keep an eye on this.


freewave serial reset installed at 9 of 11 broadbands.


See Image


Misc Problems

BROM -drifting
BFAR- spikes
FLPT- see images
DLAR.EH*- spikes
DLAR.SC* - "
MARM -see images
PARM.HN* -spikes

NM and AG Sporadic Outages - just the stations that have gone down at least the past 2nights from lack of sun


All stations were serviced. See service summary.

Three ISIS3 upgrades remain: TWAR, ARPT, and the PLAR replacement.

NODE Tasks

James has an initial mockup of a Win7 node machine with NIDAQ and adsend. Need to make sure it has all the other necessary tools and burn it in. Will try it first on mudworm then convert the BRBC machine when its pulled so that the High Peak node is Win7.


All parts in house. Will be completed immediately after the holiday break. Need a live mockup.


LNXT: ABB fades on some mornings, most likely due to atmospheric inversions. Jim noticed the side band at KLUV is at 60db instead of 80db. Or perhaps there's an issue at PWPT.


Some work remains on the NMAD tower (e.g. a couple shakey preamps). Completed, one preamp and one coax replaced.

AZTN/HDBR: NSMP reviewed the proposal to take over O&M of the bridge instrumentation. They were very complementary on the craftsmanship. But USGS is still under sequestor and is not permitted to add new stations while this remains. We need some cost estimates for replacing bad sensors and improving telemetry.


Need to review and update the power systems in 218.


Need to replace the 24V regulators at power hungry reftek sites. CCRT, and ASTN are done, PMTN got a new PS when the new media converter was installed.

CCNC is not going through fgtn and instead repeats through ASTN. The 950MHz shot would work so it could be converted to digital.

We're still repeating CPRT through GTTN. The cell modem won't work at WSNC. John will check to see if it will work at CPRT. If it does, then we can decommission the repeater at GTTN and the Seymour fire station.

A work order is in to upgrade the Athens DSL. Currently we're a 1.5Mbps download and about 350kbps upstream. The upgrade will increase to 3Mbps and 1.1Mbps.


See image
GMG.HNZ-see image

Misc Problems
BCRT.HNZ -spikes
GMG- intermittent down
ASTN.HN* - spikes
WSNC - see image

WSNC upgrade. Done.

We seem to have excessive RT147 problems. Need to follow up with ASL. Mitch will call a meeting of the FDL committee and ask for possession of the K2's. Then we can have VLF transform them into episensors. These can replace the RT147's and end the merry-go-round of replacements.

GFM needs to be moved to a more quiet and accessible site. Possible target for digital upgrade.



Internet, accessibility, and high noise inducing RF requires we move the MGNC central receive. John presented several possibilities and investigated both the PSN digitizer and the Menlo Park system. We determined that eventually we'll need to convert all nodes to the Menlo Park system so that we can obsolesce XP. But for now, we need to reconfigure North Carolina as in this figure. Better figure. This includes:
Decommission RBNC DONE
Upgrade WSNC to broadband or S-13/Reftek and cell modem. On Deck.
Rotate CCNC to repeat through fgtn and add to MTTN node. Done.
Establish new node at High Peak.
Transmit MGNC to High Peak.
Repeat TRYN through MGNC.
Transmit GFM to High Peak.
Transmit TVNC to High Peak.
Transmit SMNC to High Peak.

Node hardware at BRBC is pulled. Waiting on Morgantown cable provider to finish building upgrades. Projected completion of the upgrades is mid-February.

Need to get breakouts that connect ISIS4 to both S-13 and L4. That way we can upgrade remaining PII stations to ISIS4 first, then later go back and replace the L4's with S-13's. In progress.

BRBC Node could become the HPNC node in February/March Chris built the breakouts to connect ISIS4 to both S-13's and L-4's. We also have all the ISIS 4 AMP/VCO's and discriminators to convert the remaining 5 stations.


ASTN and PMTN would benefit from reconfiguring the basalts there to record continuous buffer in addition to the 100 sps rt data transfer.

PMTN needs a fence?


Smalley obtained 5 new antennas and 2 new choke rings. First targets for new antennas are PIGT, NWCC, MACC, and CJTR.

Need ethernet electrical isolation at all stations. MCTY, CVMS, MAIR and PIGT are top priorities.

Next time each station is visited, need to note antenna model and serno for Smalley.

Antenna problem at NWCC. Sometimes tracks all, sometimes less or 0.

AG Network

LCAR sensor installed and Leak should be fixed. Check next visit.


Chris has a pile of obsolete/unrepairable/unserviceable hardware. Need to have the committee meet to finalize inventory and decommission hardware.


NSMP Station List. All stations were converted to cell modems.


Clean out shed behind house0 to use as field storage.

The seisnet group on umdrive is at You will need your uuid and password.